# Global convergence of a modified BFGS-type method based on function information for nonconvex multiobjective optimization problems

### v1

@@ 4.2 Superlinear convergence analysis
$(1-\bar{c}\varepsilon_{k})\|\bar{s}_{k}\|\leq\|\bar{y}_{k}\|\leq(1+\bar{c}\varepsilon_{k})\|\bar{s}_{k}\|,$ (4.29)
$(1-\bar{c}\varepsilon_{k})^{2}\|\bar{s}_{k}\|^{2}-2\bar{y}_{k}^{\top}\bar{s}_{k}+\|\bar{s}_{k}\|^{2}\leq\|\bar{y}_{k}\|^{2}-2\bar{y}_{k}^{\top}\bar{s}_{k}+\|\bar{s}_{k}\|^{2}\leq\bar{c}^{2}\|\bar{s}_{k}\|^{2}\varepsilon_{k}^{2}.$
@@
$\frac{\bar{y}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}\geq 1-\bar{c}\varepsilon_{k}.$ (4.30)
Using the Assumption 4.3 (iii), and (4.30) we obtain
$\bar{a}_{k}=\frac{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}=\frac{(\bar{y}_{k}+m_{k}\nabla^{2}f_{\lambda}(x^{*})^{-1}\bar{s}_{k})^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}$ $=\frac{\bar{y}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}+m_{k}\frac{\bar{s}_{k}^{\top}\nabla^{2}f_{\lambda}(x^{*})^{-1}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}$
$\geq 1-\bar{c}\varepsilon_{k}+\frac{m_{k}}{L}$
$\geq 1-\bar{c}\varepsilon_{k}.$ (4.31)
@@
$=\frac{\|\bar{y}_{k}\|^{2}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}+2m_{k}\frac{\bar{y}_{k}^{\top}\nabla^{2}f_{\lambda}(x^{*})^{-1}\bar{s}_{k}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}+(m_{k})^{2}\frac{\bar{s}_{k}^{\top}\nabla^{2}f_{\lambda}(x^{*})^{-2}\bar{s}_{k}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}$
$\leq\frac{(1+\bar{c}\varepsilon_{k})^{2}}{1-\bar{c}\varepsilon_{k}}+2m_{k}\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\|\|y_{k}\|\|\nabla^{2}f_{\lambda}(x^{*})^{-1}\|\|\nabla^{2}f_{\lambda}(x^{*})^{\frac{1}{2}}\|\|s_{k}\|}{U\|s_{k}\|^{2}}+(m_{k})^{2}\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-2}\|\|\bar{s}_{k}\|^{2}}{U\|s_{k}\|^{2}}$
$\leq 1+\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{k}}{1-\bar{c}\varepsilon_{k}}\varepsilon_{k}+\eta_{1}m_{k}+\eta_{2}m_{k}^{2},$ (4.33)
where $\eta_{1}=2\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\|\|\nabla^{2}f_{\lambda}(x^{*})^{-1}\|\|\nabla^{2}f_{\lambda}(x^{*})^{\frac{1}{2}}\|L}{U}$, $\eta_{2}=\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-2}\|\|\nabla^{2}f_{\lambda}(x^{*})^{\frac{1}{2}}\|^{2}}{U}.$ Using (4.26) and (4.31),
$\psi(\bar{B}_{k+1})\leq\psi(\bar{B}_{k})$ $+\left(\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{k}}{1-\bar{c}\varepsilon_{k}}\varepsilon_{k}+\eta_{1}m_{k}+\eta_{2}m_{k}^{2}-\ln(1-\bar{c}\varepsilon_{k})\right)$
@@
By adding the above inequality from 0 to k, then we obtain
$0<\psi(\bar{B}_{k+1})\leq\psi(\bar{B}_{1})$ $+\sum_{j=1}^{k}\left(\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{j}}{1-\bar{c}\varepsilon_{j}}\varepsilon_{j}+\eta_{1}m_{j}+\eta_{2}m_{j}^{2}-\ln(1-\bar{c}\varepsilon_{j})\right)$
@@
From Theorem 4.4, there exists $k_{0}\in N$ such that
@@
then, $\ln(1-\bar{c}\varepsilon_{k})\geq-2\bar{c}\varepsilon_{k}.$ It follows by Assumption 4.3 (ii) that
@@
Since $-\ln(\cos^{2}\bar{\beta}_{j})>0$ ,for all $j>k_{0}$, and $h(t)=1-t+lnt\leq 0$ when $t>0$, hence, we get
$-\left[1-\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}+\ln\left(\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}\right)\right]\geq 0,$
@@
$\lim\limits_{k\to\infty}\frac{\|(B_{k}-\nabla^{2}f_{\lambda}(x^{*}))s_{k}\|^{2}}{\|s_{k}\|^{2}}$ $=\lim\limits_{k\to\infty}\frac{\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}(B_{k}-\nabla^{2}f_{\lambda}(x^{*}))\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}\bar{s}_{k}\|^{2}}{\|\nabla^{2}f_{\lambda}(x^{*})^{-\frac{1}{2}}s_{k}\|^{2}}$
@@
$=\lim\limits_{k\to\infty}\left[\frac{\bar{q}_{k}}{\cos^{2}\bar{\beta_{k}}}-2\bar{q}_{k}+1\right]$
@@
Suppose that Assumptions 4.3 hold. Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Then,
@@
(ii) $\{x_{k}\}$ converge to $x^{*}$ at a superlinear rate.
@@
The proof is similar to that in [20, Theorem4.3], we omit it here. ∎
@@ 5 Numerical experiments
For MQNMO and MFQNMO methods, we used the Frank-Wolf gradient method to solve the dual problem (3.9). For QNMO, (3.4) is solved using scipy. optimize, which is an optimization Solver. All test problems and main characteristics are listed in Table 1. Columns ‘n’ and ‘m’ indicate the number of variables and objectives, respectively. The ‘Convex’ column indicates whether the problem is convex or not. ’Y’ stands for Yes. ’N’ stands for NO. Problems are solved under box constraints ${{x}_{L}}\leq x\leq{{x}_{U}}.$ The initial points are randomly selected within these box constraints ${{x}_{L}}\leq x\leq{{x}_{U}}.$ We used the termination criterion $\left|\theta\left(x\right)\right|\leq{{10}^{-8}}$ for all test problems. The maximum number of iterations is set to 500. To obtain the stepsize that satisfies the Wolfe conditions (4.1) and (4.2), we used the Wolf line search algorithm proposed in [21]. We used $\sigma_{1}={10}^{-4}$, $\sigma_{2}=0.1$, and set $B_{0}=I_{n}$, for all $i\in[m]$. We perform 200 computations using the same initial points for all methods.
@@
In Table 2, the average values obtained from these runs are recorded to determine the number of iterations (iter), function evaluations (feval), and CPU time (time(ms)). The number of points that did not converge in 500 iterations among the 200 initial points is recorded in column ‘NF’ for MFQNMO, QNMO, and MQNMO. We denote ’Failed’ as ’F,’ which indicates that the corresponding method fails to solve the test problem under the stopping criterion. For large-scale problems, such as JOS, the QNMO method’s CPU time consumption is substantial although it requires the fewest iterations. For all nonconvex test problems in Table 1, the QNMO method exhibited poor performance and the MQNMO method surpasses the MFQNMO method only for the FF1 test problem. For certain problems in Table 1, the experimental efficiency of the MFQNMO method and the MQNMO method are equal, such as BK1, MOP1, MHHM1, and DGO1. We mentioned the possible reasons for this phenomenon in Remark 4.3. As can be seen from Table 2, the MFQNMO method is more efficient and competitive.
Figure 1 shows the Pareto frontier obtained using the MFQNMO method for the convex problem listed in Table 1. Figure 2 shows the Pareto frontier obtained using the MFQNMO method for the nonconvex problem listed in Table 1. For the chosen test problems, it can be observed that the MFQNMO method was able to satisfactorily estimate the Pareto frontiers when considering evenly distributed starting points.
(a) Lov1
(b) PNR
(c) DGO2
(d) BK1
Figure 1: The approximated Pareto frontiers generated by MFQNMO for the convex problems Lov1, PNR, DGO2, and BK1.
(a) Lov3
(b) Lov4
(c) DGO1
(d) FF1
(e) Hil
(f) KW2
(g) LDTZ
(h) MOP2
(i) SD
(j) SK1
(k) SLCDT1
Figure 2: The approximated Pareto frontiers generated by MFQNMO for the nonconvex problems Lov3, Lov4, DGO1, FF1, Hil, KW2, LTDZ, MOP2, SD, SK1, and SLDTZ1.

### v2

@@ 4.2 Superlinear convergence analysis
$(1-\bar{c}\varepsilon_{k})\|\bar{s}_{k}\|\leqslant\|\bar{y}_{k}\|\leqslant(1+\bar{c}\varepsilon_{k})\|\bar{s}_{k}\|,$ (4.30)
$(1-\bar{c}\varepsilon_{k})^{2}\|\bar{s}_{k}\|^{2}-2\bar{y}_{k}^{\top}\bar{s}_{k}+\|\bar{s}_{k}\|^{2}\leqslant\|\bar{y}_{k}\|^{2}-2\bar{y}_{k}^{\top}\bar{s}_{k}+\|\bar{s}_{k}\|^{2}\leqslant\bar{c}^{2}\|\bar{s}_{k}\|^{2}\varepsilon_{k}^{2}.$
@@
$\frac{\bar{y}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}\geqslant 1-\bar{c}\varepsilon_{k}.$ (4.31)
Using the Assumption 4.3 (ii), and (4.31) we obtain
$\bar{a}_{k}=\frac{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}=\frac{(\bar{y}_{k}+m_{k}\nabla^{2}f_{\lambda}^{k}(x^{*})^{-1}\bar{s}_{k})^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}$ $=\frac{\bar{y}_{k}^{\top}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}+m_{k}\frac{\bar{s}_{k}^{\top}\nabla^{2}f_{\lambda}^{k}(x^{*})^{-1}\bar{s}_{k}}{\|\bar{s}_{k}\|^{2}}$
$\geqslant 1-\bar{c}\varepsilon_{k}+\frac{m_{k}}{L}$
$\geqslant 1-\bar{c}\varepsilon_{k}.$ (4.32)
@@
$=\frac{\|\bar{y}_{k}\|^{2}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}+2m_{k}\frac{\bar{y}_{k}^{\top}\nabla^{2}f_{\lambda}^{k}(x^{*})^{-1}\bar{s}_{k}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}+(m_{k})^{2}\frac{\bar{s}_{k}^{\top}\nabla^{2}f_{\lambda}^{k}(x^{*})^{-2}\bar{s}_{k}}{\bar{\gamma}_{k}^{\top}\bar{s}_{k}}$
$\leqslant\frac{(1+\bar{c}\varepsilon_{k})^{2}}{1-\bar{c}\varepsilon_{k}}+2m_{k}\frac{\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}\|\|y_{k}\|\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-1}\|\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{\frac{1}{2}}\|\|s_{k}\|}{U\|s_{k}\|^{2}}+(m_{k})^{2}\frac{\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-2}\|\|\bar{s}_{k}\|^{2}}{U\|s_{k}\|^{2}}$
$\leqslant 1+\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{k}}{1-\bar{c}\varepsilon_{k}}\varepsilon_{k}+\eta_{1}m_{k}+\eta_{2}m_{k}^{2},$ (4.34)
where $\eta_{1}=2\frac{\|\ \nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}\|\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-1}\|\|\nabla^{2}f_{\lambda}(x^{*})^{\frac{1}{2}}\|L}{U}$, $\eta_{2}=\frac{\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-2}\|\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{\frac{1}{2}}\|^{2}}{U}.$ Using (4.27) and (4.32),
$\psi(\bar{B}_{k+1})\leqslant\psi(\bar{B}_{k})$ $+\left(\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{k}}{1-\bar{c}\varepsilon_{k}}\varepsilon_{k}+\eta_{1}m_{k}+\eta_{2}m_{k}^{2}-\ln(1-\bar{c}\varepsilon_{k})\right)$
@@
By adding the above inequality from $1$ to $k$, then we obtain
$0<\psi(\bar{B}_{k+1})\leqslant\psi(\bar{B}_{1})$ $+\sum_{j=1}^{k}\left(\frac{3\bar{c}+\bar{c}^{2}\varepsilon_{j}}{1-\bar{c}\varepsilon_{j}}\varepsilon_{j}+\eta_{1}m_{j}+\eta_{2}m_{j}^{2}-\ln(1-\bar{c}\varepsilon_{j})\right)$
@@
From Theorem 4.4, there exists $k_{0}\in\mathbb{N}$ such that
@@
then,
$\ln(1-\bar{c}\varepsilon_{k})\geqslant-2\bar{c}\varepsilon_{k}.$ (4.36)
We can get
$\sum\limits_{k\geqslant 1}m_{k}<\infty,$ (4.37)
from (4.24) and Theorem 4.4. It follows by (4.36) and (4.37) that
@@
Since $-\ln(\cos^{2}\bar{\beta}_{j})>0$ ,for all $j>k_{0}$, and $h(t)=1-t+lnt\leqslant 0$ when $t>0$, hence, we get
$-\left[1-\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}+\ln\left(\frac{\bar{q}_{j}}{\cos^{2}\bar{\beta_{j}}}\right)\right]\geqslant 0,$
@@
$\lim\limits_{k\to\infty}\frac{\|(B_{k}-\nabla^{2}f_{\lambda}^{k}(x^{*}))s_{k}\|^{2}}{\|s_{k}\|^{2}}$ $=\lim\limits_{k\to\infty}\frac{\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}(B_{k}-\nabla^{2}f_{\lambda}^{k}(x^{*}))\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}\bar{s}_{k}\|^{2}}{\|\nabla^{2}f_{\lambda}^{k}(x^{*})^{-\frac{1}{2}}s_{k}\|^{2}}$
@@
$=\lim\limits_{k\to\infty}\left[\frac{\bar{q}_{k}^{2}}{\cos^{2}\bar{\beta_{k}}}-2\bar{q}_{k}+1\right]$
@@
Suppose that Assumption 4.1 and Assumption 4.3 hold. Let $\{x_{k}\}$ be the sequence generated by Algorithm 1. Then,
@@
(ii) $\{x_{k}\}$ converges to $x^{*}$ at a superlinear rate.
@@
The proof is similar to that in [25, Theorem 4.3], we omit it here. ∎
@@ 5 Numerical experiments
For MQNMO and MFQNMO methods, we used the Frank-Wolf gradient method to solve the dual problem (3.9). For QNMO, (3.4) is solved using scipy.optimize, which is an optimization Solver. All test problems and main characteristics are listed in Table 1. Columns ‘n’ and ‘m’ indicate the number of variables and objectives, respectively. The ‘Convex’ column indicates whether the problem is convex or not. ‘Y’ stands for convex MOP. ‘N’ stands for nonconvex MOP. Problems are solved under box constraints ${{x}_{L}}\leqslant x\leqslant{{x}_{U}}.$ The initial points are randomly selected within these box constraints ${{x}_{L}}\leqslant x\leqslant{{x}_{U}}.$ We used the termination criterion $\left|\theta\left(x\right)\right|\leqslant{{10}^{-8}}$ for all test problems. The maximum number of iterations is set to 500. To obtain the stepsize that satisfies the Wolfe conditions (4.1) and (4.2), we used the Wolf line search algorithm proposed in [14]. We used $\sigma_{1}={10}^{-4}$, $\sigma_{2}=0.1$, and set $B_{0}=I_{n}.$ We perform 200 computations using the same initial points for all methods.
@@
In Table 2, the average values obtained from these runs are recorded to determine the number of iterations (iter), function evaluations (feval), and CPU time (time(ms)). Among the 200 initial points, the number of points that did not converge in 500 iterations is recorded in column ‘NF’ for MFQNMO, QNMO, and MQNMO. We denote ‘Failed’ as ‘F,’ which indicates that the corresponding method fails to solve the test problem under the stopping criterion. For large-scale problems, such as JOS, the QNMO method’s CPU time consumption is substantial although it requires the fewest iterations. For all nonconvex test problems in Table 1, the QNMO method exhibited poor performance and the MQNMO method surpasses the MFQNMO method only for the FF1 test problem. For certain problems in Table 1, the experimental results of the MFQNMO method and the MQNMO method are equal approximately, such as BK1, MOP1, MHHM1 and AP2. We mentioned the possible reasons for this phenomenon in Remark 4.3. As can be seen from Table 2, the MFQNMO method is more efficient and competitive.
Figure 1-3 shows the Pareto frontier obtained using the MFQNMO method for some convex problems listed in Table 1. Figure 4-10 shows the Pareto frontier obtained using the MFQNMO method for certain nonconvex problems listed in Table 1. For the chosen test problems, it can be observed that the MFQNMO method was able to satisfactorily estimate the Pareto frontiers when considering evenly distributed starting points.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 1: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the PNR problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 2: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the DGO2 problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 3: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the SD problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 4: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the FF1 problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 5: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the Hil problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 6: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the KW2 problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 7: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the SLCDT1 problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 8: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the MOP2 problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 9: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the SK1 problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 10: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the LDTZ problem.
(a) QNMO
(b) MQNMO
(c) MFQNMO
Figure 11: The approximated nondominated frontiers generated by QNMO, MQNMO, and MFQNMO for the AP3 problem.